OpenStudy (anonymous):

what is the antiderivative of 1/2x^2 -2x +6

OpenStudy (agreene):

do you mean: \[\int \frac{1}{2x^2 -2x +6 }dx\] ?

OpenStudy (anonymous):

(1/6)x^3-x^2+6x+C

OpenStudy (anonymous):

no sorry, i meant \[(1/2)x^2 -2x +6\]

OpenStudy (agreene):

In that case, cauchy has your answer.

OpenStudy (agreene):

\[\frac{x^3}{6}-x^2+6+C\] if you prefer.

OpenStudy (anonymous):

okay thanks